OpenStudy (anonymous):

HOW DO YOU SOLVE THIS ?? in 1990 tampa,florida, had a population of 280,015.from 1990 through 1996, the population increased at an average rate of about 850 per year . write a linear model for the population of tampa

OpenStudy (anonymous):

if x represents the number of years then the population will be: y = 850(x) + 280,015 check: 1990 (i.e. year 0) y = 850(0) + 280,015 = 280,015 1991 (i.e. year 1) y = 850(1) + 280,015 = 850 + 280,015 =280,865